As you zip through space in your PPS (personal propulsion suit), your pulse rate, as you count it, is 115 bpm (beats per minute). This high rate serves as objective evidence of your excitement. However, an observer on the Moon, an expert in pulse rate telemetry, measures your pulse rate as slower. In fact she detects only 0.573 times the rate you count and claims that you must be pretty calm in spite of everything that is going on. How fast are you moving with respect to the Moon?
Answer #1

Due to time dilation, the observed time is :

t_obs = t/sqrt(1-v^2/c^2)

where v = speed of the PPS with respect to moon

c = speed of light = 3*10^8 m/s

So, heat beat observed, B_obs = B*sqrt(1-v^2/c^2)

So, 0.573*115 = 115*sqrt(1-v^2/c^2)

So, 0.573^2 = (1-v^2/c^2)

So, v = 0.8196*c = 0.8196*3*10^8 = 2.46*10^8 m/s <--------answer